OpenStudy (anonymous):

solve and write interval notation for the solution set -6≤ 1/2x -4<-3

OpenStudy (anonymous):

\[-6\le(x/2)\le1\]\[-12\le(x)\le2\] [-12,2]

OpenStudy (anonymous):

wait i messed up something there, ill rewrite it

OpenStudy (anonymous):

\[-2\le(x/2)\le1\]\[-4\le(x)\le2\] so actually [-4,2]

OpenStudy (anonymous):

i forgot to add 4 to the left side

OpenStudy (anonymous):

Thank you so much i really appreciate it